find the volume of CO2 at NTP obtained on heating 4.2g of NaHCO3

Grade:11

1 Answers

Riya Rani
6 years ago
2NaHCO3 = Na2CO3 + CO2 + H2O
2 moles of NaHCO3 = 1 mole of CO2
4.2 / 84 mole of NaHCO3 = 4.2 / 168 mole of CO2
                                         = 0.025 moles of CO2
wt. = mol. wt. * moles
       = 44 * 0.025
       = 1.1gm